Use the point on the line and the slope m of the line to find three additional points through which the line passes.
(3,-6), m = 0

Answer by mananth(16946)

y=mx +c
-6 =0+C
c=-6
y=-6 is the equation of the line
(4,-6)
(5,-6)
(6,-6)